THE PERIMETER OF A RECTANGLE IS 147METERS. THE LENGHT IS 15 METERS LESS THAN TIMES THE WIDTH. FIND THE DIMENSSIONS. THANK YOU
Answer by ewatrrr(24785)

 
Hi,

P = L + w = 73.5m
(??w-15m) + w = 73.5 the 'times' part did not come thru